OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x))

OpenStudy (goformit100):

The Warmest Welcome to Open Study.

OpenStudy (anonymous):

merci

jimthompson5910 (jim_thompson5910):

f(x) = 9x - 2 f(g(x)) = 9( g(x) ) - 2 ... Replace every x with g(x) f(g(x)) = 9( -x + 3 ) - 2 ... On the right side, plug in g(x) = -x + 3 I'll let you simplify that

OpenStudy (anonymous):

I got 8x=22, simplifying it would turn it to a fraction, what then? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

simplify 9( -x + 3 ) - 2 and tell me what you get

OpenStudy (anonymous):

-9x + 25

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

f(g(x)) = -9x+25 is the answer

OpenStudy (anonymous):

oh wow I did so much unnecessary extra work, thanks bro for the help

jimthompson5910 (jim_thompson5910):

you're welcome